# Bidirectional Prototype-Reward co-Evolution for Test-Time Adaptation of Vision-Language Models

### v1

@@ 3.3 Prototype-Reward Interactive Evolution
where $R_{i}$ is the comprehensive quality score derived from our multi-dimensional reward mechanism, and $\tau$ is a temperature parameter that controls the sharpness of the weight distribution.
Method ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average OOD Average
CLIP-ResNet-50 [23] 58.16 21.83 51.41 56.15 33.37 44.18 40.69
Ensemble 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp [43] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT [25] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [8] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [14] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS [27] 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS [41] 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE [37] 63.41 30.15 56.72 63.72 40.03 50.81 47.66
BPRE (Ours) 63.77 30.77 56.97 63.94 39.91 51.07 47.90
$\Delta$0.11 -0.29 +0.25 +0.22 -0.12 +0.26 +0.24
CLIP-ViT-B/16 [23] 66.73 47.87 60.86 73.98 46.09 59.11 57.20
Ensemble 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp [43] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT [25] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [8] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [14] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS [27] 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS [41] 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE [37] 71.91 59.63 65.44 80.40 52.26 65.93 64.43
BPRE (Ours) 72.22 60.67 65.50 80.68 53.45 66.50 65.08
$\Delta$0.03 +0.56 +0.07 +0.28 +0.25 +0.57 +0.65
Table 1: Robustness Evaluation under Natural Distribution Shifts. We report comparative top-1 accuracy (%) across methods, utilizing CLIP’s ResNet-50 and ViT-B/16 architectures as visual encoders. Bold values indicate optimal performance.
@@
Overall, the dual promotion mechanism creates a dynamic feedback loop between reward estimation and prototype evolution, facilitating their mutual enhancement. The mechanism operates through two complementary processes: (1) high-fidelity prototypes serve as anchor points for precise reward assessment, and (2) refined reward signals guide targeted prototype updates. As illustrated in Fig. 4, increasing the number of iterative samples results in tighter clustering of class-specific features, with a corresponding rise in reward scores from the Multi-Dimensional Quality-Aware Reward Module. This iterative interaction progressively enhances the framework’s ability to model fine-grained visual-semantic relationships and adapt to intra-class variations. By co-evolving prototypes and rewards, the mechanism improves robustness against noisy data and fosters interpretable cross-modal feature learning, ultimately producing more discriminative and generalizable representations for cross-modal recognition.
4 Experiments
We assess the BPRE framework through a thorough experimental protocol, focusing on three key dimensions: distributional robustness, cross-dataset generalization across 15 benchmarks, and inference efficiency. Additionally, systematic ablation studies are conducted to validate the core design principles guiding our architectural decisions.
CLIP-ResNet50 [23] 15.66 85.88 55.70 40.37 23.69 61.75 73.97 83.57 58.80 58.84 55.82
Ensemble 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 56.63
CoOp [43] 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
TPT [25] 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT [8] 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA [14] 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE [37] 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
BPRE (Ours) 22.05 91.24 59.71 52.96 44.74 69.79 77.59 87.35 64.88 63.13 63.34
$\Delta$ +2.25 +0.41 -1.00 +2.78 +2.63 +1.05 -1.62 +1.17 +0.65 -1.05 +1.41
CLIP-ViT-B/16 23.67 93.35 65.48 44.27 42.01 67.44 83.65 88.25 62.59 65.13 63.58
Ensemble 23.22 93.55 66.11 45.04 50.42 66.99 82.86 86.92 65.63 65.16 64.59
CoOp [43] 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
TPT [25] 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT [8] 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA [14] 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE [37] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
BPRE(Ours) 29.76 95.17 67.73 55.32 55.94 76.86 86.36 92.48 70.24 71.13 70.10
$\Delta$ +0.81 +0.36 +0.42 +1.12 -2.06 +1.79 -0.87 +1.34 +0.17 +0.47 +0.70
Table 2: Cross-Dataset Generalization Analysis. We evaluate top-1 accuracy (%) across methods using both CLIP backbones (ResNet-50 and ViT-B/16). Bold entries denote superior performance.
4.1 Datasets
@@
4.2 Implementation Details
We adopt PyTorch as the deep learning framework and utilize CLIP with ResNet-50 and ViT-B/16 backbones as the visual encoders. Following TPT, each test image is augmented into 64 views through random resized cropping (63 augmented + 1 original). The Dynamic Prototype Evolution employs a momentum-based update mechanism (momentum=0.9) with an Multi-Dimensional Quality-Aware Reward Module. For the prototype evolution function, we set temperature $\tau=0.01$. The cache mechanism maintains a queue of size 3 per class, with a normalized entropy threshold of 0.1 for prototype updates. During testing, the reward calculator undergoes a 1000-step warmup period with a minimum reward threshold of 0.1, while the prototype manager performs weighted feature averaging every 10 iterations based on computed rewards. All experiments are conducted on either a single NVIDIA RTX 4090 or A100 GPU. For feature space visualization, we utilize t-SNE with perplexity 30 to project the prototype features into 2D space.
4.3 Comparison with State-of-the-art
To validate the effectiveness of our approach, we conduct comprehensive comparisons against several state-of-the-art test-time adaptation methods for CLIP [23]. As a fundamental baseline, we include CLIP’s zero-shot performance using both single prompt (”a photo of a CLASS”) and prompt ensemble strategies. Our comparative analysis encompasses various methodological categories. In the prompt optimization domain, we consider TPT [25], which leverages self-entropy minimization across augmented views, and its advanced variant DiffTPT [8], which incorporates diffusion-based augmentations. For cache-based approaches, we evaluate TDA [14], which employs training-free adapters with positive-negative cache construction, and DMN-ZS [41], which features backpropagation-free dynamic memory aggregation. We further compare against prototype-focused methods, including TPS [27], which efficiently learns per-class prototype shifts, and DPE [37], which implements dual-modal prototype enhancement during testing. To ensure fair evaluation, all baseline results are directly cited from their original publications.
@@ 4.3 Comparison with State-of-the-art
Table 3: Efficiency comparison on ImageNet [5]. We report the testing time, the achieved accuracy, and the performance gains compared to zero-shot CLIP.
Robustness to Distribution Shifts. In Tab.  1, we present comprehensive evaluations on ImageNet and its four challenging out-of-distribution variants. Our BPRE demonstrates consistent improvements over existing approaches across both ResNet-50 and ViT-B/16 backbones. While baseline CLIP [23] shows reasonable in-domain performance, its effectiveness deteriorates under distribution shifts. Though traditional enhancement methods like Ensemble and CoOp [43] offer moderate improvements, their gains remain limited. Among test-time adaptation methods, BPRE achieves state-of-the-art performance across both ResNet-50 and ViT-B/16 architectures. Specifically, with ResNet-50, BPRE surpasses prompt-tuning based approaches (TPT [25] by 3.81% and DiffTPT [8] by 2.36%), cache-based method TDA [14] by 1.49%, and recent methods like DMN-ZS [41] and DPE [37] by 1.10% and 0.26% respectively. Similar performance gains are observed with ViT-B/16, where BPRE outperforms TPT by 4.06%, DiffTPT by 4.22%, TDA by 1.49%, and demonstrates consistent improvements over DMN-ZS (+1.01%) and DPE (+0.57%). These comprehensive results across different architectures validate BPRE’s superior adaptation capability and robustness. These improvements, though modest in absolute terms, are significant given the already strong performance of recent methods. The consistent gains across different architectures and datasets validate the effectiveness of our approach in handling distribution shifts without requiring additional training data.
Cross-Datasets Generalization. Our extensive evaluation in Tab.  2 demonstrates BPRE’s superior performance across 10 diverse datasets using two prominent backbone architectures. With ResNet-50, BPRE achieves significant improvements over the CLIP baseline (+7.52%) and consistently outperforms state-of-the-art methods, including TPT (+5.68%), DiffTPT (+3.49%), TDA (+2.31%), and DPE (+1.41%). The performance gains are particularly notable on challenging fine-grained datasets, with substantial improvements on Aircraft (+2.25%) and DTD (+2.78%). For ViT-B/16, BPRE establishes a new state-of-the-art benchmark with 70.10% average accuracy, demonstrating consistent advantages over leading competitors DPE (+0.70%) and TDA (+2.57%). The method exhibits remarkable domain adaptability, showing significant improvements on specialized datasets like Flower (+1.79%) and Pets (+1.34%). These comprehensive results validate BPRE’s effectiveness as a lightweight test-time adaptation approach that requires neither additional training data nor complex optimization strategies.
Efficiency Comparison. Tab.  3 compares the efficiency of our method with other VLM test-time adaptation approaches on 5,000 ImageNet test samples using a 24GB NVIDIA RTX 4090 GPU. Our model requires only 27 minutes more than the latest DPE [37] model while achieving a 0.36% improvement in prediction accuracy.
4.4 Ablation Studies
Effectiveness of the Multi-Dimensional Quality-Aware Reward Module. To evaluate the contribution of each component in our proposed Multi-Dimensional Quality-Aware Reward Module, we performed an extensive analysis on the ImageNet dataset. As presented in Tab. 4, the removal of the $R_{sim}$, $R_{conf}$, and $R_{div}$ components individually results in a marked degradation in model performance. Specifically, when removing $R_{sim}$, the model’s ability to capture semantic similarities is compromised; without $R_{conf}$, the confidence-based quality assessment is impaired; and the absence of $R_{div}$ reduces the model’s capacity to maintain sample diversity. These findings empirically validate our hypothesis that multi-dimensional evaluation metrics are essential for robust sample quality assessment.
Ablation studies on different values of queue size $M$. Tab. 5 illustrates our systematic investigation of queue size $M$ sensitivity on the ImageNet dataset. Our empirical results demonstrate that $M=3$ yields optimal performance, establishing a critical equilibrium between two competing factors: insufficient intra-class diversity at lower $M$ values and the introduction of detrimental low-confidence samples at higher $M$ values. This finding underscores the importance of appropriate queue size selection in maintaining robust model performance.
# $R_{sim}$ $R_{conf}$ $R_{div}$ Acc.
1 63.48
2 63.56
3 63.59
4 63.52
5 63.61
6 63.65
7 63.64
8 63.77
Table 4: Ablation study of different components of the Multi-Dimensional Quality-Aware Reward Module.
Ablation studies on different values of $R_{min}$. A systematic evaluation of the minimum reward threshold $R_{min}$ was conducted on the ImageNet dataset using a ResNet-50 backbone, as presented in Tab.  6. The experimental results conclusively demonstrate that $R_{min}=0.1$ maximizes model effectiveness. This optimal threshold value plays a crucial role in maintaining the delicate balance between model robustness and discriminative power, highlighting the significance of precise parameter tuning in our proposed framework.
@@ 4.4 Ablation Studies
Acc. 63.31 63.52 63.77 63.56 63.72
Table 5: Ablation studies on different values of $M$.
More Sensitivity Analyses of Hyper-Parameters. We conducted a comprehensive hyperparameter sensitivity analysis on the ImageNet dataset using a ResNet-50 backbone, with results visualized in Fig.  5. Fig.  5(a) reveals that the model achieves peak performance when both temperature parameters are set to $\alpha=4$ and $\beta=4$. As illustrated in Fig.  5(b), the optimal balance among reward components is achieved with $\lambda_{sim}=0.6$ and $\lambda_{conf}=0.2$ (with $\lambda_{div}=0.2$ to satisfy the constraint $\lambda_{sim}+\lambda_{conf}+\lambda_{div}=1$).
Values of $R_{min}$ 0.0 0.01 0.05 0.1 0.2 0.3 0.4
ImageNet Acc. 62.76 63.23 63.54 63.77 63.68 63.65 63.59
Table 6: Ablation studies on different values of $R_{min}$.

### v2

@@ III-C2 Quality-Aware Visual Prototype Evolution
Here, $R_{final}$ is the comprehensive quality score from our multi-dimensional reward mechanism, and $\tau$ controls the weight distribution’s sharpness.
@@
Overall, our dual promotion mechanism establishes a dynamic feedback loop between reward estimation and prototype evolution. This operates through two complementary processes: high-fidelity prototypes enable precise reward assessment, while refined reward signals guide targeted prototype updates. As shown in Fig. 4, increasing iterative samples leads to tighter class-specific feature clustering and higher reward scores. This co-evolution progressively enhances the framework’s ability to model visual-semantic relationships and adapt to intra-class variations, improving robustness against noise and fostering interpretable cross-modal feature learning for more discriminative representations.
TABLE II: Robustness Evaluation under Natural Distribution Shifts. We report top-1 accuracy (%) across methods using CLIP’s ResNet-50 and ViT-B/16 backbones. Bold and underlined values indicate the best and second-best performance, respectively.
Method Source Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets SUN397 UCF101 Average
CLIP-ResNet50 [6] ICML’21 15.66 85.88 55.70 40.37 23.69 61.75 73.97 83.57 58.80 58.84 55.82
CoOp [26] IJCV’22 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
CoCoOp [27] CVPR’22 14.61 87.38 56.22 38.53 28.73 65.57 76.20 88.39 59.61 57.10 57.23
Tip-Adapter [7] ECCV’22 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 56.63
TPT [13] NeurIPS’22 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT [14] ICCV’22 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA [15] CVPR’24 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE [16] NeurIPS’24 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
BPRE Ours 22.05 91.24 59.71 52.96 44.74 69.79 77.59 87.35 64.88 63.13 63.34
CLIP-ViT-B/16 ICML’21 23.67 93.35 65.48 44.27 42.01 67.44 83.65 88.25 62.59 65.13 63.58
CoOp [26] IJCV’22 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
CoCoOp [27] CVPR’22 22.29 93.79 64.90 45.45 39.23 70.85 83.97 90.46 66.89 68.44 64.63
Tip-Adapter [7] ECCV’22 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 56.63
TPT [13] NeurIPS’22 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT [14] ICCV’23 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA [15] CVPR’24 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE [16] NeurIPS’24 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
BPRE Ours 29.76 95.17 67.73 55.32 55.94 76.86 86.36 92.48 70.24 71.13 70.10
IV Experiments
We evaluate BPRE using a comprehensive evaluation protocol that emphasizes distributional robustness and cross-dataset generalization. Furthermore, systematic ablation studies are carried out to validate the efficacy of our proposed design principles.
IV-A Datasets
@@
We implement our approach using PyTorch with CLIP (ResNet-50 and ViT-B/16 architecture). Following TPT, each test image was augmented into 64 views (63 augmented + 1 original). The dynamic prototype evolution utilized a momentum-based update mechanism (momentum = 0.9) with our Multi-dimensional Quality-aware Reward Module. For prototype evolution, we set temperature $\tau=0.01$. Our cache maintains 3 samples per class with a normalized entropy threshold of 0.1. During testing, the reward calculator underwent a 1000-step warmup with a minimum threshold of 0.1, while prototype updates occur every 10 iterations through weighted feature averaging. Experiments were conducted on a single NVIDIA RTX 4090, and we adopted t-SNE (perplexity 30) to project prototype features into a 2D space.
TABLE III: Efficiency comparison on ImageNet [39]. We reported testing time, achieved accuracy, and performance gains compared to zero-shot CLIP model.
@@
IV-B Comparison with SOTA Methods
To validate our approach, we compared several state-of-the-art TTA methods for CLIP [6]. We included the zero-shot performance of CLIP using both single-prompt and prompt-ensemble strategies. Our analysis encompassed prompt optimization methods (TPT [13], including DiffTPT [14]), cache-based approaches (TDA [15], DMN-ZS [35]), and prototype-focused methods (TPS [38], DPE [16]). It is noted that all comparison results are directly cited from their original publications.
IV-B1 Robustness to Distribution Shifts
To evaluate the robustness of our method under natural distribution shifts, we conducted experiments on ImageNet and four challenging out-of-distribution (OOD) variants. As summarized in Tab. I, BPRE consistently outperformed existing approaches across both ResNet-50 and ViT-B/16 backbones, demonstrating strong robustness to domain shifts. Specifically, while baseline CLIP [6] achieved reasonable in-domain accuracy on ImageNet, its performance degraded sharply when evaluated on OOD datasets such as ImageNet-A and ImageNet-Sketch, highlighting its vulnerability to domain shifts. Several methods, such as CoOp [26] or CoCoOp [27], provided only moderate improvements, indicating that prompt learning alone is insufficient for robust test-time adaptation, especially in the absence of target domain labels. Among recent test-time adaptation strategies, BPRE achieved superior performance on both architectures. For the ResNet-50 backbone, BPRE surpasses prompt-tuning approaches (TPT [13] by 3.81%, DiffTPT [14] by 2.36%), cache-based TDA [15] by 1.49%, and prototype-based methods (DMN-ZS [35] and DPE [16] by 1.10% and 0.26%, respectively). A similar trend was observed with the ViT-B/16 backbone, where BPRE outperforms TPT by 4.06%, DiffTPT by 4.22%, TDA by 1.49%, DMN-ZS by 1.01%, and DPE by 0.57%. Notably, BPRE achieved the largest margins on the most challenging OOD datasets, such as ImageNet-A and ImageNet-Sketch, where the distribution gap is the greatest. These improvements were particularly meaningful given the strong performance of existing baselines and the inherent challenge of TTA without access to source data or target labels. The consistent gains across OOD benchmarks indicate that the bidirectional prototype-reward co-evolution mechanism is highly effective in identifying high-quality features and dynamically refining class prototypes, thereby achieving better semantic alignment with the target domain even under severe distribution shifts. Integrating similarity, confidence, and diversity within its multi-dimensional reward module, BPRE mitigated over-reliance on text priors and reduced misclassification, resulting in cross-modal misalignment. This positive feedback loop between rewards and prototypes enables progressive knowledge refinement, resulting in not only higher accuracy but also more stable and reliable performance across diverse scenarios. Overall, these results validated the robustness and generalizability of BPRE when domain shift is prevalent.
TABLE IV: Comparison with DPE on the SigLIP across 10 datasets.
DPE (SigLIP) 10.44 62.39 14.76 61.88 38.58 72.68 26.18 79.34 66.96 55.30 48.85
Ours (SigLIP) 13.82 64.26 15.83 62.58 38.40 73.04 41.83 79.37 67.35 52.74 50.92
IV-B2 Cross-Datasets Generalization
To assess the cross-dataset generalization capability of BPRE, we conducted experiments on 10 diverse benchmarks using two backbone architectures. As shown in Tab. II, BPRE consistently achieved superior performance compared to existing methods. BPRE with ResNet-50 achieves significant improvements over the CLIP model (+7.52%) and consistently outperforms state-of-the-art methods, including TPT (+5.68%), DiffTPT (+3.49%), TDA (+2.31%), and DPE (+1.41%). The performance gains are particularly notable on challenging fine-grained datasets, with substantial improvements on Aircraft (+2.25%) and DTD (+2.78%), where subtle visual distinctions between classes require more precise prototype representations. For ViT-B/16, BPRE established a new SOTA benchmark with 70.10% average accuracy, demonstrating consistent advantages over leading competitors DPE (+0.70%) and TDA (+2.57%). The method exhibits remarkable domain adaptability, showing significant improvements on specialized datasets such as Flower (+1.79%) and Pets (+1.34%). BPRE maintained strong performance on several datasets with varying numbers of classes, image resolutions, and visual features. Unlike the OOD experiments on ImageNet variants that primarily assess robustness to broad distribution shifts, fine-grained benchmarks highlight the need to distinguish subtle and localized visual differences between similar categories. The consistent superiority of BPRE in these tasks stems from its bidirectional prototype-reward co-evolution mechanism, which iteratively refines class prototypes using multi-dimensional feature quality assessment, integrating similarity, confidence, and diversity to select features that are both representative and discriminative. Our method is particularly effective where class boundaries are ambiguous and other methods may lack sufficient granularity. Furthermore, BPRE features a lightweight, training-free design that enables efficient adaptation without requiring additional data or complex optimization. Overall, these results demonstrated our effectiveness and generalizability for the unique challenges of fine-grained visual tasks.
TABLE V: Ablation studies of different components of the MQRM on the Food and ImageNet datasets.
# Components Food ImageNet
$R_{\text{sim}}$ $R_{\text{conf}}$ $R_{\text{div}}$ Acc. (%) Acc. (%)
1 76.12 63.56
2 75.97 63.59
3 74.21 63.52
4 76.68 63.61
5 75.78 63.65
6 74.42 63.64
7 77.59 63.77
IV-B3 Efficiency Comparison.
To evaluate the efficiency of BPRE, we compared its efficiency and accuracy with state-of-the-art test-time adaptation (TTA) methods on 50,000 ImageNet test samples. As shown in Tab. III, while TPT and DiffTPT achieved modest gains (+0.93% and +0.99%), their high computational costs (>10h and >20h) make them impractical for practical use. In contrast, BPRE achieves an optimal balance between efficiency and performance, requiring only slightly more computation than DPE (3h 04min vs. 2h 37min) while delivering higher accuracy (+0.36%). These results demonstrate that our bidirectional co-evolution mechanism enhances model generalization without significant computational overhead, making BPRE well-suited for test-time adaptation scenarios where both accuracy and efficiency are crucial.
IV-C Ablation Studies
IV-C1 Robustness Analysis on Different Architectures
To further assess the robustness and generalizability of our method, we evaluated it on the SigLIP architecture and compared it with DPE on 10 diverse datasets, as shown in Tab. IV. Our approach consistently surpassed DPE (SigLIP) on nearly all benchmarks, with notable improvements on datasets such as Aircraft (13.82% vs. 10.44%), Food101 (41.83% vs. 26.18%), and Caltech (64.26% vs. 62.39%). Our method achieved a 2.07% higher accuracy (50.92% vs. 48.85%) on average. These results demonstrated that our approach maintained strong performance across a range of visual domains and dataset characteristics, including both fine-grained and large-scale recognition tasks. The consistent gains also indicated that our method was compatible with different backbone architectures, such as SigLIP, and was not limited by specific model designs. These findings further demonstrate the robustness and versatility of our approach in adapting to various architectural settings and data distributions.
TABLE VI: Ablation study on different values of $M$.
@@ IV-C1 Robustness Analysis on Different Architectures
ImageNet Acc. (%) 63.31 63.52 63.77 63.56 63.72
Aircraft Acc. (%) 19.56 21.77 22.05 22.01 20.42
IV-C2 Effectiveness of the MQRM
We conducted an ablation study on Food101 [37] to evaluate each component in our Multi-dimensional Quality-aware Reward Module. As shown in Tab. V, removing any reward component consistently degrades performance. Among single-component experiments, $R_{sim}$ (76.12%) provides the strongest signal, followed by $R_{conf}$ (75.97%) and $R_{div}$ (74.21%). Dual-component configurations show improved results, with $R_{sim}$+$R_{conf}$ achieving the highest accuracy (76.68%). The full implementation delivered superior performance (77.59%), surpassing the best dual-component configuration by 0.91%. This progressive improvement demonstrated the complementary nature of our reward components, where each addresses a distinct aspect of feature quality assessment. The observed results were attributed to the specific roles each reward component plays in the adaptation process. The similarity reward ($R_{sim}$) ensures that selected features are closely aligned with class prototypes, providing a strong foundation for semantic consistency. The confidence reward ($R_{conf}$) favors features with high prediction certainty, which helps reduce the influence of noisy or ambiguous samples. The diversity reward ($R_{div}$) encourages the selection of a broader range of informative features, preventing overfitting to a narrow subset of the data. When used in combination, these components not only reinforced strengths but also compensated for their individual limitations. For instance, while $R_{sim}$ alone may overlook uncertain yet informative features, adding $R_{conf}$ and $R_{div}$ ensures that the model remains robust to uncertainty and captures a richer set of discriminative cues. The best performance achieved by the full reward module highlights the importance of a holistic approach to feature quality assessment, enabling more effective and stable prototype refinement during test-time adaptation.
TABLE VII: Ablation study on different values of $R_{\text{min}}$.
$R_{\text{min}}$ 0.0 0.05 0.1 0.2 0.4
ImageNet Acc. (%) 62.76 63.54 63.77 63.68 63.59
Food Acc. (%) 76.31 76.89 77.59 77.37 77.03
IV-C3 Ablation studies on different values of queue size $M$